Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 07:54:17 on June 28, 2010
Entry number 328891
Followups:
keyword=Owl shift summary: June 28, 2010
Swing Shift Summary: June 26, 2010
Shift Crew
- Pete Markowitz
- Aidan Kelleher
- Eli Piazetsky
Event Summary
- 00:00 Shift started with no beam, Bogdan working on RHRS detectors.
MCC thought to be several hours from beam.
- 00:25 Markowitz, Kelleher go downstairs to work on wiring up S2m.
Vince comes in to fix clock for scalers. 00:09 Out of hall, MCC ready to
deliver beam.
- 01:09 First beam in hall. MCC tuning. Put in Tantalum target but
see no rate(?) T1 and T2 are about 10% greater than cosmic with 0.5uA.
Head scratchings, magnet checks, etc.
- 02:47 Move to thick carbon target. See T2~3 KHz with 1 uA.
- 03:00 Access to fix TDC spectra. Bogdan reverses earlier cable
change and TDC spectra are no longer empty.
- Q1 current oscillations return. Tried trick from last shift (turn
off regulation, set current to proper value, turn on regulation) but that
no longer works. Since we are not doing optics, small oscillations may
be tolerable now but day shift should ask a tech to look at this.
- 03:49 Septum back on, running thick carbon and spectra no longer
empty.
- 03:59 Move to Tantalum. Still no rate. Maybe the Ta target is
gone? Rates stay about cosmic rates (T1~1KHz, T2~140 Hz, T5=T1~1KHz).
- 04:21 Move to Pb-D1. Healthy rate (T1~16 KHz, T2~8.9KHz) with 1
uA.
- 04:45 Septum off for access. Bogdan will put gas cerenkov into T5
trigger. T1 is S2m, T2 is S0, and T5 is T1 currently. Sergei has a nice
plot showing positrons in the lead glass detector.
- 05:55 Out of hall, septum back on. T5 should now have the gas
cerenkov in coincidence with T2, the S0 trigger (T1 is out of the T5
trigger now). Ask for 1 uA on the Pb-D1 target.
- 06:03 Ask MCC to take away beam while we analyze run 1282, to keep
radiation in hall low. Run 1282 should have T5 as positron rate.
- 06:10 Beam off rates are 0.5 Hz for T5, 400 Hz for T2. With 1 uA on
the Pb-D1, we saw T2~14KHz and T5~532Hz. I.e., after subtracting
cosmics, the pion rates are about 7.3KHz and the positron rates about
about 532 Hz, or a ratio of ~24:1 if this was a good trigger.
- 06:20 Move the thick carbon target to check ratio of rates there.
Run 1283, 1uA on the thick carbon. Rates are 2.3KHz for T2 (with ~400 Hz
being cosmics), 14Hz for T5 giving a ratio of 140:1.
- 06:35 HV were changed. Move back to Pb-D1. Run 1284 is with the new
gas cerenkov HV.
- 06:40 Move to thick carbon and start run 1285.
- 07:07 Change gas cerenkov HV again. Move target to PbD1, start run
1286
- 07:15 More HV changes. Start moving to 12C thick, but reconsider and
stay on Pb-D1. Start run 1287.
- 07:30 More HV changes. Run 1288.
- 07:40 Final HV changes. Run 1289. At end, ask for hall to go to
controlled access for day shift.
Issues that need sumpin'
- End of run statements give essentially zero for BCM average current
values. Maybe something is not calibrated? Runs have also been very
short too.
- Q1 current/field oscillates. Swing shift had thought this was fixed
but apparently not.
- Tantalum target foil is probably gone.
- Spot is not working.